# Towards the New XAI: A Hypothesis-Driven Approach to Decision Support Using Evidence

### v3

@@ 4.5 Conduct
where $u(h)$ is the uncertainty level of hypothesis $h$ given the probabilistic output is $p(h)$. To ensure there was a clear difference between high and low uncertainty, we select instances with low uncertainty by choosing instances entropy less than 0.3, and high uncertainty by choosing instances with entropy greater than 0.7. Participants did not know the certainty nor how many test instances were correct/incorrect. Each participant was paid a minimum of £4 for their time, plus a bonus of £2 if they could answer at least 9 out of 12 questions correctly. Participants were also given a plain language statement, and consent form and did a training phase with 3 example questions before answering the 12 test questions.
For RQ4 the text is a response to “Can you please explain why you selected this option?". We analysed a total of 12 (questions) $\times$ 95 (participants) = 1140 responses. The final analysis includes 1,031 responses after removing 109 responses due to poor quality. Each response is assigned to at least one category (or code): Using feature values or Using evidence. We then perform a simple deductive analysis by reading each response and assign the relevant codes. We explain each code as follows: (1) Using Feature Values: participants rely on the feature values and their background knowledge to make the final decision without using the model evidence; and (2) Using Evidence: Participants rely on the evidence provided by the model and possibly their background knowledge to make the final decision.
We chose these two codes based on the idea of machine explanation and human intuition [8, 9]. Specifically, using evidence refers to using the machine explanation and therefore, making use of the model evidence to support decision-making. On the other hand, using feature values is relevant to using people’s intuitions of the task based on the input feature values. Therefore, using the qualitative analysis, we explore how people use the model evidence and their intuitions in the three decision-making paradigms.
@@ 4.6 Participants
Using the power analysis for F-test for one factor ANOVA and assuming the power of 0.8 and significant alpha of 0.05, we found that a sample size of 300 participants in three groups guarantees a small effect size of 0.2. In total, we recruited $N=302$ participants on Prolific, distributed into three conditions: 102 participants in C1, 99 participants in C2 and 101 participants in C3. Participants are selected from the United States, United Kingdom, and Australia and must be fluent in English. Gender-wise, 192 were women, 103 were men, 4 self-specified their gender and 3 declined to state their gender. Age-wise, 94 participants were between Ages 18 and 29, 91 were between Ages 30 and 39, 44 were between Ages 40 and 49, and 73 were over Age 50.
@@
5 Experiment Results
In this section, we show the results of two experiments. In the first experiment, we explore whether hypothesis-driven can improve task efficiency, task performance and reduce reliance compared to recommendation-driven and AI explanation only. In the second experiment, we understand how participants used our hypothesis-driven approach differently compared to the other two baselines. The full statistics and screenshots from our user study can be found in the supplementary material [25].
@@ 5 Experiment Results
Time (minutes) $\mathbf{17.92}$  $\pm$ $\mathbf{9.34}$ 18.62  $\pm$ 11.26 18.09  $\pm$ 9.26 $p=0.808,r=0.020$ $p=0.993,r=0.001$ $p=0.892,r=0.011$
Brier score 0.29  $\pm$ 0.07 0.30  $\pm$ 0.07 $\mathbf{0.27}$  $\pm$ $\mathbf{0.06}$ $p=0.566,r=0.047$ $\mathbf{p=0.016,r=0.196}$ $\mathbf{p=0.003,r=0.245}$
Over-reliance (%) 73.86  $\pm$ 20.91 54.21  $\pm$ 22.51 $\mathbf{53.30}$  $\pm$ $\mathbf{22.73}$ $\mathbf{p<0.001,r=0.450}$ $\mathbf{p<0.001,r=0.449}$ $p=0.946,r=0.006$
Under-reliance (%) $\mathbf{17.81}$  $\pm$ $\mathbf{20.35}$ 41.25  $\pm$ 27.18 24.42  $\pm$ 18.19 $\mathbf{p<0.001,r=0.523}$ $\mathbf{p=0.001,r=0.246}$ $\mathbf{p<0.001,r=0.387}$
Table 1: Results of the human experiment. R: Recommendation-driven, O: AI-explanation-only, H: Hypothesis-driven. Winners/significances are highlighted.
5.1 Experiment 1: Quantitative Results
We performed a Shapiro-Wilks test to check the data normality and we found that our data was not normally distributed ($p<0.05$). Therefore, we apply non-parametric Kruskal-Wallis test. We then perform post-hoc Mann-Witney U test to do pairwise comparisons. The results are visualised in Table 1 and Table 1. The significant differences between two conditions are highlighted in italic red in the figures where $p<0.05$.
@@ 5.1 Experiment 1: Quantitative Results
We evaluate participants’ decision-making performance by using the Brier score. A lower Brier score indicates better decision accuracy. As seen in Table 1, participants in the hypothesis-driven condition ($M=0.267$, $SD=0.063$) have a lower Brier score than the other two approaches (C1: ($M=0.290$, $SD=0.071$), C2: ($M=0.295$, $SD=0.073$)). We accept H2a/b. Therefore, hypothesis-driven helps participants be confident when they make a correct decision, and be less confident when they make a wrong decision.
@@
In Table 1, hypothesis-driven ($M=53.30$, $SD=22.73$) reduced over-reliance significantly compared to recommendation-driven ($M=73.86$, $SD=20.91$) ($p=1.5\times 10^{-8}$, $r=0.449$). We accept H3a. Moreover, AI-explanation-only ($M=54.21$, $SD=22.51$) also reduces over-reliance compared to the recommendation-driven approach ($p=1.6\times 10^{-8}$, $r=0.450$).
@@
In Table 1, hypothesis-driven ($M=24.42$, $SD=18.19$) significantly reduced under-reliance compared to AI-explanation-only ($M=41.25$, $SD=27.18$) ($p=1.09\times 10^{-6}$, $r=0.387$). We accept H3b. This is not surprising because we expect that participants in the AI-explanation-only condition are the most likely to ‘under-rely’ due to being unable to compare evidence across hypotheses nor see a recommendation. Recommendation-driven ($M=17.81$, $SD=20.35$) has the least under-reliance value because participants were given recommendations.
@@ 5.2 Experiment 2: Qualitative Results
In Figure 1 and 2, we illustrate the number of times that participants used feature values and evidence to make their decisions based on the text analysis.
For the recommendation-driven paradigm, participants use the feature values to confirm whether the decision aid’s prediction and explanation are reliable or not. If participants think the feature values do not match the evidence explanation, they will go with the feature values to make the final decision. Some examples that the study participants in the recommendation-driven condition go against the decision aid’s prediction:
@@
Ignoring evidence is, of course, a good strategy if the decision maker believes that the evidence is wrong. However, recommendation-driven does not help participants to be aware of the high uncertainty among multiple predictions. This limitation is mitigated by the hypothesis-driven paradigm.
For the AI explanation only paradigm, participants often rely on the feature values and not on the evidence explanation to make a decision. This is not surprising because participants can find it difficult to interpret the evidence without seeing the label that the evidence is referred to. We attribute this to the cognitive effort to link evidence to hypotheses, leading to participants ignoring evidence and relying on input feature values to make their decisions. This is a noteworthy limitation of AI-explanation-only as it makes people overlook the explanation if the link to the evidence is unclear. In the study by Gajos and Mamykina [15], the link from feature attributions to the task solution is more straightforward than in our study, which may explain the divergence of results.
Participants more often use the evidence to make a decision in the hypothesis-driven approach than in recommendation-driven or AI explanation only. This shows participants took advantage of the model evidence. In the two baseline conditions, participants tended to ignore evidence seemingly due to the inability to interpret it, which means they will fail to take advantage of the underlying model. In Figure 1, there are only two exceptions at Q6 and Q11 where the evidence is not the most used in the hypothesis-driven condition.
@@
    I was choosing between high and medium. Quality of construction, age and location are the most important features. When it was at high, these were all positive. Kitchen quality and fireplaces were negative, but these are not as important.” – Q0

### v4

@@ 4.5 Conduct
where $u(h)$ is the uncertainty level of hypothesis $h$ given the probabilistic output is $p(h)$. To ensure there was a clear difference between high and low uncertainty, we selected instances with low uncertainty by choosing instances with entropy less than 0.3, and high uncertainty by choosing instances with entropy greater than 0.7. Participants did not know the certainty nor how many test instances were correct/incorrect. Each participant was paid a minimum of £4 for their time, plus a bonus of £2 if they could answer at least 9 out of 12 questions correctly. Participants were also given a plain language statement, and consent form and did a training phase with 3 example questions before answering the 12 test questions.
For RQ4 the text is a response to “Can you please explain why you selected this option?". We analysed a total of 12 (questions) $\times$ 95 (participants) = 1140 responses. The final analysis includes 1,031 responses after removing 109 responses due to poor quality. Each response is assigned to at least one category (or code): Using feature values or Using evidence. We then perform a simple deductive analysis by reading each response and assigning the relevant codes. We explain each code as follows: (1) Using Feature Values: participants rely on the feature values and their background knowledge to make the final decision without using the model evidence; and (2) Using Evidence: Participants rely on the evidence provided by the model and possibly their background knowledge to make the final decision.
We chose these two codes based on the idea of machine explanation and human intuition [14, 15]. Specifically, using evidence refers to using the machine explanation and therefore, making use of the model evidence to support decision-making. On the other hand, using feature values is relevant to using people’s intuitions of the task based on the input feature values. Therefore, using the qualitative analysis, we explore how people use the model evidence and their intuitions in the three decision-making approaches.
@@ 4.6 Participants
Using the power analysis for the F-test for one factor ANOVA and assuming the power of 0.8 and significant alpha of 0.05, we found that a sample size of 300 participants in three groups guarantees a small effect size of 0.2. In total, we recruited $N=302$ participants on Prolific, distributed into three conditions: 102 participants in C1, 99 participants in C2 and 101 participants in C3. Participants are selected from the United States, United Kingdom, and Australia and must be fluent in English. Gender-wise, 192 were women, 103 were men, 4 self-specified their gender and 3 declined to state their gender. Age-wise, 94 participants were between Ages 18 and 29, 91 were between Ages 30 and 39, 44 were between Ages 40 and 49, and 73 were over Age 50. All collected data were included in the analysis, as there was no evidence indicating that any participants provided poor-quality data.
@@
4.7 Experiment 1: Quantitative Results
Table 1: Results of the human experiment. R: Recommendation-driven, O: AI-explanation-only, H: Hypothesis-driven. O-R: Over-Reliance. U-R: Under-Reliance. Lower is better in all measures ($\downarrow$). Winners/significances are indicated in bold.
@@ 4.7 Experiment 1: Quantitative Results
Time (min) $\mathbf{17.92}$  $\pm$ $\mathbf{9.34}$ 18.62  $\pm$ 11.26 18.09  $\pm$ 9.26 $p=0.808,r=0.02$ $p=0.993,r=0.00$ $p=0.892,r=0.01$
Brier score 0.29  $\pm$ 0.07 0.30  $\pm$ 0.07 $\mathbf{0.27}$  $\pm$ $\mathbf{0.06}$ $p=0.566,r=0.05$ $\mathbf{p=0.016,r=0.20}$ $\mathbf{p=0.003,r=0.25}$
O-R (%) 73.86  $\pm$ 20.91 54.21  $\pm$ 22.51 $\mathbf{53.30}$  $\pm$ $\mathbf{22.73}$ $\mathbf{p<0.001,r=0.45}$ $\mathbf{p<0.001,r=0.45}$ $p=0.946,r=0.01$
U-R (%) $\mathbf{17.81}$  $\pm$ $\mathbf{20.35}$ 41.25  $\pm$ 27.18 24.42  $\pm$ 18.19 $\mathbf{p<0.001,r=0.52}$ $\mathbf{p=0.001,r=0.25}$ $\mathbf{p<0.001,r=0.39}$
In this section, we show the results of experiment 1, in which we explore whether hypothesis-driven can improve task efficiency and task performance, and reduce reliance compared to recommendation-driven and AI-explanation-only.
We performed a Shapiro-Wilks test to check the data normality and we found that our data was not normally distributed ($p<0.05$). Therefore, we apply non-parametric Kruskal-Wallis test. We then perform post-hoc Mann-Witney U test to do pairwise comparisons. The results are shown in Table 1. The significant differences between the two conditions are indicated in bold where $p<0.05$.
@@
We evaluate participants’ decision-making performance by using the Brier score. A lower Brier score indicates better decision accuracy. As seen in Table 1, participants in the hypothesis-driven condition ($M=0.267$, $SD=0.063$) have a lower Brier score than the other two approaches (C1: ($M=0.290$, $SD=0.071$), C2: ($M=0.295$, $SD=0.073$)). We accept H2a/b with small effect sizes. Therefore, hypothesis-driven helps participants be confident when they make a correct decision, and be less confident when they make a wrong decision.
@@
In Table 1, hypothesis-driven ($M=53.30$, $SD=22.73$) reduced over-reliance significantly compared to recommendation-driven ($M=73.86$, $SD=20.91$) ($p=1.5\times 10^{-8}$, $r=0.449$). We accept H3a with a medium effect size of $0.45$. Moreover, AI-explanation-only ($M=54.21$, $SD=22.51$) also reduces over-reliance compared to the recommendation-driven approach ($p=1.6\times 10^{-8}$, $r=0.450$).
@@
In Table 1, hypothesis-driven ($M=24.42$, $SD=18.19$) significantly reduced under-reliance compared to AI-explanation-only ($M=41.25$, $SD=27.18$) ($p=1.09\times 10^{-6}$, $r=0.387$). We accept H3b with a medium effect size of $0.39$. This is not surprising because we expect that participants in the AI-explanation-only condition are the most likely to ‘under-rely’ due to being unable to compare evidence across hypotheses nor see a recommendation. Recommendation-driven ($M=17.81$, $SD=20.35$) has the least under-reliance value because participants were given recommendations.
Subjective Questionnaire
We also asked participants to answer a subjective questionnaire after the task. The questionnaire consists of twelve questions to measure In control, Preference, Mental demand, System complexity and Trust. Details of the questions are provided in online supplementary materials.
Table 2: Results of the subjective questionnaire. R: Recommendation-driven, O: AI-explanation-only, H: Hypothesis-driven. Values are ranged from 0 (disagree strongly) to 10 (agree strongly). Winners/significances are indicated in bold.
R O H R vs. O R vs. H O vs. H
In control ($\uparrow$) $\mathbf{7.22}$  $\pm$ $\mathbf{2.44}$ 5.73  $\pm$ 2.83 7.01  $\pm$ 2.46 $\mathbf{p<0.001,r=0.32}$ $p=0.496,r=0.06$ $\mathbf{p<0.001,r=0.27}$
Preference ($\uparrow$) $\mathbf{5.51}$  $\pm$ $\mathbf{2.67}$ 4.05  $\pm$ 2.79 5.41  $\pm$ 3.03 $\mathbf{p<0.001,r=0.30}$ $p=0.930,r=0.01$ $\mathbf{p=0.002,r=0.26}$
Mental demand ($\downarrow$) $\mathbf{3.96}$  $\pm$ $\mathbf{2.64}$ 5.41  $\pm$ 2.68 4.54  $\pm$ 2.80 $\mathbf{p<0.001,r=0.30}$ $p=0.168,r=0.11$ $\mathbf{p=0.028,r=0.18}$
Complexity ($\downarrow$) $\mathbf{4.38}$  $\pm$ $\mathbf{2.53}$ 5.65  $\pm$ 2.91 4.42  $\pm$ 2.95 $\mathbf{p=0.001,r=0.26}$ $p=0.899,r=0.01$ $\mathbf{p=0.003,r=0.24}$
Trust ($\uparrow$) 4.92  $\pm$ 1.69 3.62  $\pm$ 1.95 $\mathbf{5.16}$  $\pm$ $\mathbf{1.98}$ $\mathbf{p<0.001,d=0.72}$ $p=0.352,d=0.13$ $\mathbf{p<0.001,d=0.79}$
In Table 2, AI-explanation-only is significantly worse than the other conditions in all facets. Moreover, there is no statistical difference between recommendation-driven and hypothesis-driven. The reason is that we conduct between-subject experiments so each participant has access to only one condition and they do not have another condition to compare to. If we run within-subject experiments to measure subjective questions in the future, participants can compare different decision-making approaches and evaluate which one they prefer the most.
Notably, the hypothesis-driven condition did not result in the highest mental demand. This is because the AI-explanation-only condition was more difficult to interpret. In the AI-explanation-only condition, participants did not know which hypothesis the evidence was referring to, making it challenging to make sense of the information. Moreover, when participants relied on their prior knowledge, they could check their hypothesis and assess the evidence to confirm it, which is easier than trying to understand a machine recommendation when it differs.
@@ 4.8 Experiment 2: Qualitative Results
In Figures 3 and 4, we illustrate the number of times that participants used feature values and evidence to make their decisions based on the text analysis. The questions in this section (Q0-Q11) refer to the 12 house instances used in the experiment described in Section 4.5. Participant comments are attributed to the question number, not the participant ID.
For the recommendation-driven approach, participants use the feature values to confirm whether the decision aid’s prediction and explanation are reliable or not. If participants think the feature values do not match the evidence explanation, they will go with the feature values to make the final decision. Some examples that the study participants in the recommendation-driven condition go against the decision aid’s prediction:
@@
Ignoring evidence is, of course, a good strategy if the decision-maker believes that the evidence is wrong. However, recommendation-driven does not help participants to be aware of the high uncertainty among multiple predictions. This limitation is mitigated by the hypothesis-driven approach.
For the AI-explanation-only approach, participants often rely on the feature values and not on the evidence explanation to make a decision. This is not surprising because participants can find it difficult to interpret the evidence without seeing the label that the evidence is referred to. We attribute this to the cognitive effort to link evidence to hypotheses, leading to participants ignoring evidence and relying on input feature values to make their decisions. This is a noteworthy limitation of AI-explanation-only as it makes people overlook the explanation if the link to the evidence is unclear. In the study by Gajos and Mamykina [23], the link from feature attributions to the task solution is more straightforward than in our study, which may explain the divergence of results.
Participants more often use the evidence to make a decision in the hypothesis-driven approach than in recommendation-driven or AI-explanation-only. This shows participants took advantage of the model evidence. In the two baseline conditions, participants tended to ignore evidence seemingly due to the inability to interpret it, which means they will fail to take advantage of the underlying model. In Figure 3, there are only two exceptions at Q6 and Q11 where the evidence is not the most used in the hypothesis-driven condition.
@@
    I was choosing between high and medium. Quality of construction, age and location are the most important features. When it was high, these were all positive. Kitchen quality and fireplaces were negative, but these are not as important.” – Q0